OpenStudy (anonymous):

the altitude of a triangle is increasing at a rate of 1 cm/min while the area in increasing at a rate of 2 cm^2/min. At what rate is the base of the triangle changing when the altitude is 10 cm and the area is 100cm^2?

OpenStudy (anonymous):

I suppose \[ A = \frac 12 bh \]And thus by power rule \[ \frac {dA}{dt}=\frac 12\frac{db}{dt}h+\frac 12 b\frac {dh}{dt} \]

hero (hero):

Area of Triangle is \[A = \frac{bh}{2}\] If we take the derivative, we get: \[A'(t) = \frac{b'h + bh'}{2}\] After inputting the given values we have: \[100 = \frac{b(10)}{2}\] \[b = 20\] And \[2 = \frac{b'(10) + 20(1)}{2}\] Solving for \(b'\) we get: \[b' = \frac{4 - 20}{10} = -\frac{8}{5}\] Therefore the rate at which the base of the triangle is changing is -8/5 cm/min